OpenStudy (anonymous):

fully factor x^4-5x^2+4

OpenStudy (anonymous):

\[x^4-5x^2+4\]

Directrix (directrix):

What are two numbers that multiply to 4 and add to -5?

OpenStudy (anonymous):

but dont we only use that method when it's a quadratic ? :S

Directrix (directrix):

You can think of this problem (quartic) in the same way as you do the quadratic as you begin factoring. So, what are the two numbers that multiply to 4 and add to -5?

OpenStudy (anonymous):

-4,-1

Directrix (directrix):

Then, x^4-5x^2+4 = (x^2 - 4) (x^2 -1) Now, factor (x^2 - 4) and next factor (x^2 -1). You should end up with four linear factors for the factorization of x^4-5x^2+4. Post what you get here.

OpenStudy (anonymous):

(x+2)(x-2)(x+1)(x-1)

OpenStudy (anonymous):

difference of squares

Directrix (directrix):

Correct. Factor: x^4 - 1.

OpenStudy (anonymous):

\[(x^2-1)(x^2+1)\]

OpenStudy (anonymous):

\[x^2+1\] ^ that can't be factored further, right?

Directrix (directrix):

x^2 + 1 cannot be further factored over the set of Reals but x^2 -1 can. x^4 - 1 = (x^2 +1)(x^2-1) and then ....

OpenStudy (anonymous):

and then (x-1)(x+1)

Directrix (directrix):

Yes. On your paper, you would write the complete factorization: x^4 - 1 = (x^2 +1)(x^2-1) = (x^2 +1)(x +1) (x - 1)

OpenStudy (anonymous):

thanks for your help, i have another factoring question that i need help on, can you please help ?

Directrix (directrix):

@burhan101 Yes. Post it here.